OpenStudy (anonymous):

log a(x)= -1 , log a (y)= 2 , and log a(z)= 5 then what does log a(x^2/yz^4) = ?

OpenStudy (anonymous):

log (xy) = log (x) + log(y) expand the last expression and add up the terms.

OpenStudy (anonymous):

i don't understand what your saying

OpenStudy (anonymous):

do you understand what log (xy) = log (x) + log(y) means?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, you can extend that to log(x/y) = log(x) - log(y) Now use these tricks over and over again on log a(x^2/yz^4)

OpenStudy (ash2326):

log x=-1 log y=-2 log z=5 i'm not writing the base, it's assumed to be a now we have to find log(x^2/(yz^4)) log(x^2)-log(y)-log(z^4) 2log(x)-log(y)-4log(z) 2*(-1)-2-4*5 -2-2-20 -24 ans